THE PURE ROTATIONAL SPECTRUM OF SrSH (X̃^{2}A)

\underline{D. T. HALFEN}, A. J. APPONI, L. M. ZIURYS, University of Arizona.

The pure rotational spectrum of SrSH (X̃^{2}A) in the vibrational ground state has been measured using direct absorption mm/sub-mm spectroscopy. The radical was produced by the reaction of vaporized Sr and H_{2}S gas in the presence of a DC discharge. Ten a-type dipole allowed transitions with K_{a} asymmetry components up to K_{a} = 6 have been recorded in the range of 350-405 GHz. Preliminary rotational constants (MHz) have been determined and are listed below. A full analysis of the spectrum, as well as a structure, will be presented.

A = 283504, B = 2877.287, C = 2845.508, (ϵ_{bb} + ϵ_{cc})/2 = 57.2